Write the slope of 7x+5y=35, give the slope of the line, give the y-intercept and graph the line
1
Expert's answer
2020-06-28T17:11:55-0400

Solution:

Linear function.

To find the line slope and y-intercept, we need to bring the function to the form:


"y=m\u00d7x+b."


Where m - slope of the line, b - y-intercept.


"7\u00d7x+5\u00d7y=35;\n5\u00d7y=-7\u00d7x+35;""y=(-7\u00d7x+35)\u00f75=-7\u00f75\u00d7(x+5)=-1.4\u00d7(x+5)=-1.4\u00d7x+7."

"y=-1.4\u00d7x+7."

The slope of the line m = -1.4, y-intercept b = 7.


A linear function is constructed

using two points.


"x=0;7\u00d70+5\u00d7y=35;5\u00d7y=35;y=35\u00f75=7."

First point A(0,7).


"y=0;7\u00d7x+5\u00d70=35;7\u00d7x=35;x=35\u00f77=5."

Second point B(5,0).


Answer: m = -1.4; b = 7; A(0,7); B(5,0).


Graph of function.